# latexText page

### v1

@@ 2.3 Backtesting dynamic ES contributions
We introduce a setting for estimating dynamic ES contributions. Let $\{\bm{X}_{t}\}_{t\in\mathbb{N}}$, $\bm{X}_{t}=(X_{1,t},\dots,X_{d,t})$, be a series of losses of interest, which is adapted to the filtration ${\cal G}=\{{\cal G}_{t}\}_{t\in\mathbb{N}}$. For $S_{t}=X_{1,t}+\cdots+X_{d,t}$, let $VaR_{t}$ and $ES_{t}$ be VaR and ES of $S_{t}|{\cal G}_{t-1}$, respectively, with the prescribed confidence level $\alpha$. For $j=1,\dots,d$, the ES contribution of $X_{j,t}$ to $S_{t}$ given ${\cal G}_{t-1}$ is denoted by $ESC_{j,t}$.
@@
We introduce a generic notation $\bm{\varrho}_{t}$ for the pair $(ESC_{j,t},VaR_{t})$ for a fixed $j\in\{1,\dots,d\}$, or the $(d+1)$-tuple $(ESC_{1,t},\dots,ESC_{d,t},VaR_{t})$. Let $\{\hat{\bm{\varrho}}_{t}\}_{t\in\mathbb{N}}$ and $\{\hat{\bm{\varrho}}_{t}^{\ast}\}_{t\in\mathbb{N}}$ be two series of predictions of $\{\bm{\varrho}_{t}\}_{t\in\mathbb{N}}$, which are $\mathcal{G}$-predictable. We assume that ~$\tilde{\mathcal{F}}_{\text{c}}^{1}(\mathbb{R}^{d+1})$ in (3) contains all distributions of $(\bm{X}_{t},S_{t})$ and $(\bm{X}_{t},S_{t})|\mathcal{G}_{t-1}$ almost surely. The multi-objective elicitability and identifiability presented in Section 2.2 enable the conduct of comparative backtests and traditional backtests (calibration tests) based on the following statistics:
$\mathbf{\bar{S}}(\hat{\bm{\varrho}})$ $=\frac{1}{|\mathbb{T}_{\text{out}}|}\sum_{t\in\mathbb{T}_{\text{out}}}\mathbf{S}(\hat{\bm{\varrho}}_{t},\bm{X}_{t}),$
$\mathbf{\bar{V}}(\hat{\bm{\varrho}})$ $=\frac{1}{|\mathbb{T}_{\text{out}}|}\sum_{t\in\mathbb{T}_{\text{out}}}\mathbf{V}(\hat{\bm{\varrho}}_{t},\bm{X}_{t}),$
with $\mathbf{\bar{S}}(\hat{\bm{\varrho}}^{\ast})$ and $\mathbf{\bar{V}}(\hat{\bm{\varrho}}^{\ast})$ defined analogously, where $\mathbf{S}$ is a multi-objective scoring function for $\bm{\varrho}$, $\mathbf{V}$ is an identification function for $\bm{\varrho}$, and $\mathbb{T}_{\text{out}}$ is the out-of-sample period $\mathbb{T}_{\text{out}}$ such that we forecast and backtest $\{\bm{\varrho}_{t}\}_{t\in\mathbb{T}_{\text{out}}}$. In practice, we divide a given finite sample period $\mathbb{T}=\{1,\dots,n+T\}$ into the in-sample period $\mathbb{T}_{\text{in}}=\{1,\dots,n\}$ and the out-of-sample period $\mathbb{T}_{\text{out}}=\{n+1,\dots,n+T\}$. For each $t\in\mathbb{T}_{\text{out}}$, we forecast $\bm{\varrho}_{t}$ at $t-1$ based on the past $n$ observations of $\{\bm{X}_{t-s}\}_{s\in\mathbb{T}_{\text{in}}}$.
Comparative backtests of ES contributions concern whether some order between $\mathbf{\bar{S}}(\hat{\bm{\varrho}})$ and $\mathbf{\bar{S}}(\hat{\bm{\varrho}}^{\ast})$ is statistically supported, and one-step and two-step approaches are proposed due to multi-objective elicitability with respect to the lexicographical order; see Section 5 of Fissler and Hoga (2023) for details. A key property of scoring functions regarding the comparative backtests is order sensitivity, which addresses comparability of two misspecified forecasts. We defer some technical discussion on this property in Section S1.2. To summarize it briefly, the average scores of the two misspecified forecasts of ES contributions are ordered when (1) the corresponding total VaRs are ordered (in the component-wise sense in Definition 3.1 of Fissler and Ziegel (2019)), or (2) the misspecified total VaRs are equal (in this case, $\operatorname{S}_{1}^{\text{ESC}},\dots,\operatorname{S}_{d}^{\text{ESC}}$ elicit biased ES contributions), the two misspecified ES contributions are ordered in the component-wise sense, and they are not between the true and biased ES contributions. Regarding Case (2), misspecification of the total VaR produces intervals such that multiple forecasts of ES contributions belonging there have inconsistent orders of average scores.
In traditional backtests of ES contributions, we seek statistical evidence on the signs of $\mathbf{\bar{V}}(\hat{\bm{\varrho}})$ and $\mathbf{\bar{V}}(\hat{\bm{\varrho}}^{\ast})$; see Section 2.2 of Nolde and Ziegel (2017) for details. It is useful to associate the sign of $\mathbf{\bar{V}}$ with over- and under-estimations of ES contributions. This is because under-estimation of risk functionals is considered to be more problematic than over-estimation from a regulatory viewpoint. With the identification functions in Proposition 2, we have, for $\bm{X}\in\mathcal{F}$, that $\mathbb{E}[\operatorname{V}^{\text{VaR}}(v,S)]\leq 0$ if and only if $\operatorname{VaR}_{\alpha}(S)\leq v$; i.e., the forecast $v$ over-estimates the total VaR. Moreover, under the correct specification $v=\operatorname{VaR}_{\alpha}(S)$ and for $j\in\{1,\dots,d\}$, a forecast $m_{j}\in\mathbb{R}$ of the $j$th ES contribution is over-estimated; i.e., $\operatorname{ESC}_{\alpha}(X_{j},S)\leq m_{j}$, if and only if $\mathbb{E}[\operatorname{V}_{j}^{\text{ESC}}((m_{j},v),\bm{X})]\leq 0$.
@@ 3 Robust forecast evaluation of ES contributions
To conduct a comparative backtest as presented in Section 2.3, we select a (multi-objective) scoring function from the class of functions characterized in Proposition 1. This is also required when we estimate models via score minimization, which we will consider in Section 4. Such a dependence of the risk measurement procedure on the choice of scoring function complicates the fair evaluation of forecasts. Patton (2020) also shows that forecast rankings can be sensitive to the choice of scoring function.
@@
is called an elementary scoring function for VaR. For competing forecasts $\{\widehat{VaR}_{t}^{(\ell)}\}_{t\in\mathbb{T}_{\text{out}}}$, $\ell=1,\dots,L$, of VaRs, a Murphy diagram displays the curves:
$\eta\mapsto\frac{1}{|\mathbb{T}_{\text{out}}|}\sum_{t\in\mathbb{T}_{\text{out}}}\operatorname{S}_{\eta}^{\text{VaR}}\left(\widehat{VaR}_{t}^{(\ell)},s_{t}\right),\quad\ell=1,\dots,L,$
against $\eta$. The diagram is negatively oriented in the sense that forecasts with lower curves are evaluated to be more accurate.
The next proposition provides mixture representations of the scoring functions for ES contributions presented in Proposition 1.
@@
    The scoring function $\operatorname{S}^{\text{ESC}}$ in Proposition 1 (S2), with $\phi_{j}=\cdots=\phi_{d}=:\phi$, admits the mixture representation:
@@
The elementary scoring function (6) can be interpreted as a degree of regret in the following setting. Suppose that the $j$th branch of a company has fixed capital $\eta$ to cover a future loss, whose point forecast is denoted by $m_{j}$, incurred at this branch under the specific situation when the company incurs a loss greater than $v$. If $m_{j}\leq\eta$, then the branch expects that the initial capital can cover a future loss. In this case, since an excess loss $(x_{j}-\eta)$ is incurred when the realized loss $x_{j}$ exceeds $\eta$ and $s>v$ occurs, the amount $\operatorname{S}_{j,\eta}^{\text{ESC}}((m_{j},v),\bm{x})=\mathbbm{1}{\{s>v\}}(x_{j}-\eta)_{+}$ can be understood as a degree of regret against the initial expectation. If $\eta<m_{j}$, then the branch expects that the initial capital does not cover a future loss, and thus some risk treatment can be conducted. In this case, the branch realizes the squandered opportunity of capital reduction $(\eta-x_{j})_{+}$ when $s>v$ occurs, and thus the degree of regret against the initial expectation can be measured by $\operatorname{S}_{j,\eta}^{\text{ESC}}((m_{j},v),\bm{x})=\mathbbm{1}{\{s>v\}}(\eta-x_{j})_{+}$. Provided $v=\operatorname{VaR}_{\alpha}(S)$, the true $j$th ES contribution is the optimal $m_{j}$ minimizing the expected degree of regret.
Since $\phi_{j}$ and $\phi$ in (M1) and (M2) above are strictly convex, the corresponding measures $H_{j}$ and $H$ assign positive mass to any finite interval on $\mathbb{R}$. Note that $\mathrm{d}H(x)=2\mathrm{d}x$ for the square loss, and $\operatorname{S}_{j,\eta}^{\text{ESC}}$ arises from $\operatorname{S}_{j}^{\text{ESC}}$ in Proposition 1 (S1) by taking $\phi_{j}(x)=(x-\eta)_{+}$, although this function is not strictly convex.
Based on the mixture representations (5) and (7), a Murphy diagram of ES contributions can be drawn analogously to the case of VaR. For $l\in\{1,\dots,L\}$, let:
$\{\hat{\bm{\varrho}}_{t}^{(\ell)}\}_{t\in\mathbb{T}_{\text{out}}},\quad\text{where}\quad\hat{\bm{\varrho}}_{t}=(\widehat{ESC}_{1,t}^{(\ell)},\dots,\widehat{ESC}_{d,t}^{(\ell)},\widehat{VaR}_{t}^{(\ell)}),$
be a series of predictions in the setting of Section 2.3. The Murphy diagram of the $j$th ES contribution displays the curves.
@@
The Murphy diagram of the $d$-tuple of ES contributions analogously exhibits $\eta\mapsto\bar{\operatorname{S}}_{\eta}^{\text{ESC}}(l):=\sum_{j=1}^{d}\bar{\operatorname{S}}_{j,\eta}^{\text{ESC}}(l)$ for $\ell=1,\dots,L$. We obtain these curves and their differences on the whole real line by evaluating them on a finite set. This is because $\bar{\operatorname{S}}_{j,\eta}^{\text{ESC}}(l)$ and $\bar{\operatorname{S}}_{\eta}^{\text{ESC}}(l)$ are piecewise linear functions of $\eta$ with all kinks and jump points contained in $\mathcal{D}_{j}^{(l)}=\{x_{j,t},\widehat{ESC}_{j,t}^{(\ell)},t\in\mathbb{T}_{\text{out}}\}$ and $\mathcal{D}^{(l)}=\bigcup_{j=1}^{d}\mathcal{D}_{j}^{(l)}$, respectively. Moreover, $\bar{\operatorname{S}}_{j,\eta}^{\text{ESC}}(l)$ and $\bar{\operatorname{S}}_{\eta}^{\text{ESC}}(l)$ vanish outside of $[\min(\mathcal{D}_{j}^{(l)}),\max(\mathcal{D}_{j}^{(l)})]$ and $[\min(\mathcal{D}^{(l)}),\max(\mathcal{D}^{(l)})]$, respectively.
We note that Murphy diagrams of ES contributions depend on the corresponding forecast of total VaR. Therefore, as naturally induced from multi-objective elicitability, a completely fair evaluation among multiple forecasts of ES contributions can be conducted when their corresponding forecasts of total VaR are all equal. In such a case, more advanced analyses, such as forecast dominance tests proposed by Ziegel et al. (2020), are available.
@@ 4 Proposed models for estimating dynamic ES contributions
Pprevious sections have shown that clear and rigid comparison among forecasting models of ES contributions is accessible among those with equal total VaRs. In this section we exploit this feature and propose a new model based on compositinal regression.
@@
Despite the appeals of the top-down approach in forecasting ES contributions, special attention is required in this approach so that the empirical counterpart of the full allocation property $\sum_{j=1}^{d}\widehat{ESC}_{j,t}=\widehat{ES}_{t}$ is satisfied. To handle this constraint, we propose a new model for estimating dynamic ES contributions. In what follows, we describe our proposed procedure for the case of one-step ahead forecast, where we use the losses $\{\bm{X}_{t}\}_{t\in\mathbb{T}_{\text{in}}}$, $\mathbb{T}_{\text{in}}=\{1,\dots,n\}$, to forecast ES contributions at time $n+1$.
The proposed model consists of two stages, where the dynamics of total ES are first estimated in combination with that of total VaR, and then ES contributions are estimated from the proportion of the total ES to each component of risk. In the first stage, we estimate the dynamics of $(ES_{t},VaR_{t})$ for $t\in\mathbb{T}_{\text{in}}$ and then forecast $(VaR_{n+1},ES_{n+1})$ based on this estimated dynamics. For this purpose, various models are proposed in the literature; see, for example, Patton et al. (2019)Taylor (2019) and Taylor (2022). Joint elicitability of the pair of risk measures $(\operatorname{VaR}_{\alpha},\operatorname{ES}_{\alpha})$ enables us to estimate the joint dynamics of this pair by score minimization; see, for example, Taylor (2019). In this paper we do not specify any specific model in this stage and instead select the best model based on the currently available information and candidate models. Denote by $(\widehat{VaR}_{t},\widehat{ES}_{t})_{t\in\mathbb{T}}$, $\mathbb{T}=\{1,\dots,n+1\}$, the dynamics forecasted in this stage.
In the second stage, we allocate $\widehat{ES}_{t}$ to each component of risk with the weight vector $\bm{w}_{t}=(w_{1,t},\dots,w_{d,t})\in[0,1]^{d}$, $\sum_{j=1}^{d}w_{j,t}=1$, such that $ESC_{j,t}=w_{j,t}\times ES_{t}$. For all $j=1,\dots,d$ and $t\in\mathbb{T}$, we assume that $10<w_{j,t}<$ 1, which is implied by the natural assumption that $0<ESC_{j,t}<ES_{t}$. Consequently, the weight vector $\bm{w}_{t}$ lies in $\mathcal{S}_{d}=\left\{\bm{w}\in(0,1)^{d}:\sum_{j=1}^{d}w_{j}=1\right\}$. Data on $\mathcal{S}_{d}$ are called compositional data, and statistical modeling of such data has been extensively studied in the literature; see Aitchison (1982); Aitchison and J. Egozcue (2005); Pawlowsky-Glahn and Buccianti (2011); and references therein. A common approach for modeling compositional data is first transforming the data to an unconstrained space to eliminate the sum constraint and then constructing a model on the unconstrained space.
One of the most prominent examples of such a transform is the isometric logratio transformation (ilr, Egozcue et al. 2003) defined by:
where $V\in\mathbb{R}^{d\times(d-1)}$ is a given matrix such that $V^{\top}V=I_{d-1}$, $VV^{\top}=I_{d}-\frac{1}{d}\mathbf{1}_{d}\mathbf{1}_{d}^{\top}$ and $V^{\top}\mathbf{1}_{d}=\mathbf{0}_{d-1}$. Matrix $V$ is called the contrast matrix. The inverse map of $\operatorname{ilr}$ is given by:
@@
With this choice of $V$, $\operatorname{ilr}_{k}(\bm{w})$, $k\in\{1,\dots,d-1\}$, can be interpreted as a balance between the $k$th weight and the sum of the weights over the assets $k+1,\dots,d$. Note that 00 and $1$ cannot be included in the compositional data for the above transformations to be well-defined. This well-known limitation of ilr does not cause a significant problem in our analysis since 00 or $1$ in the allocated weights corresponds to the practically irrelevant case when no capital or full capital is allocated to one asset, respectively.

### v2

@@ 2.3 Backtesting dynamic ES contributions
We introduce a setting for estimating dynamic ESCs. Let $\{\bm{X}_{t}\}_{t\in\mathbb{N}}$, $\bm{X}_{t}=(X_{1,t},\dots,X_{d,t})^{\top}$, be a series of losses of interest, which is adapted to the filtration ${\cal G}=\{{\cal G}_{t}\}_{t\in\mathbb{N}}$. For $S_{t}=X_{1,t}+\cdots+X_{d,t}$, let $VaR_{t}$ and $ES_{t}$ be VaR and ES of $S_{t}|{\cal G}_{t-1}$, respectively, with the prescribed confidence level $\alpha$. For $j=1,\dots,d$, the $j$th ESC of $\bm{X}_{t}$ given ${\cal G}_{t-1}$ is denoted by $ESC_{j,t}$.
@@
We introduce a generic notation $\bm{\varrho}_{t}$ for the pair $(ESC_{j,t},VaR_{t})$ for a fixed $j\in\{1,\dots,d\}$, or the $(d+1)$-tuple $(ESC_{1,t},\dots,ESC_{d,t},VaR_{t})$. Let $\{\hat{\bm{\varrho}}_{t}\}_{t\in\mathbb{N}}$ and $\{\hat{\bm{\varrho}}_{t}^{\ast}\}_{t\in\mathbb{N}}$ be two series of ($\mathcal{G}$-predictable) forecasts of $\{\bm{\varrho}_{t}\}_{t\in\mathbb{N}}$. We assume that ~$\tilde{\mathcal{F}}_{\text{c}}^{1}(\mathbb{R}^{d+1})$ in (3) contains all distributions of $(\bm{X}_{t},S_{t})$ and $(\bm{X}_{t},S_{t})|\mathcal{G}_{t-1}$ almost surely. The multi-objective elicitability and identifiability presented in Section 2.2 enable the conduct of comparative backtests and traditional backtests (calibration tests) based on the statistics $\mathbf{\bar{S}}(\hat{\bm{\varrho}})=(1/|\mathbb{T}_{\text{out}}|)\sum_{t\in\mathbb{T}_{\text{out}}}\mathbf{S}(\hat{\bm{\varrho}}_{t},\bm{X}_{t})$ and $\mathbf{\bar{V}}(\hat{\bm{\varrho}})=(1/|\mathbb{T}_{\text{out}}|)\sum_{t\in\mathbb{T}_{\text{out}}}\mathbf{V}(\hat{\bm{\varrho}}_{t},\bm{X}_{t})$, respectively, with $\mathbf{\bar{S}}(\hat{\bm{\varrho}}^{\ast})$ and $\mathbf{\bar{V}}(\hat{\bm{\varrho}}^{\ast})$ defined analogously, where $\mathbf{S}$ is a multi-objective scoring function for $\bm{\varrho}$, $\mathbf{V}$ is an identification function for $\bm{\varrho}$, and $\mathbb{T}_{\text{out}}$ is the out-of-sample period $\mathbb{T}_{\text{out}}$ such that we forecast and backtest $\{\bm{\varrho}_{t}\}_{t\in\mathbb{T}_{\text{out}}}$. In practice, we divide a given finite sample period $\mathbb{T}=\{1,\dots,n+T\}$ into the in-sample period $\mathbb{T}_{\text{in}}=\{1,\dots,n\}$ and the out-of-sample period $\mathbb{T}_{\text{out}}=\{n+1,\dots,n+T\}$. For each $t\in\mathbb{T}_{\text{out}}$, we forecast $\bm{\varrho}_{t}$ at $t-1$ based on the past $n$ observations of $\{\bm{X}_{t-s}\}_{s\in\mathbb{T}_{\text{in}}}$.
Comparative backtests of ESCs concern whether some order between $\mathbf{\bar{S}}(\hat{\bm{\varrho}})$ and $\mathbf{\bar{S}}(\hat{\bm{\varrho}}^{\ast})$ is statistically supported, and one-step and two-step approaches are proposed due to multi-objective elicitability with respect to the lexicographical order; see Section 5 of Fissler and Hoga (2024) for details. In traditional backtests of ESCs, we seek statistical evidence on the signs of $\mathbf{\bar{V}}(\hat{\bm{\varrho}})$ and $\mathbf{\bar{V}}(\hat{\bm{\varrho}}^{\ast})$. We refer the reader to Section 2.2 of Nolde and Ziegel (2017) for details.
@@ 3 Robust forecast evaluation of ES contributions
To conduct a comparative backtest as presented in Section 2.3, we select a (multi-objective) scoring function from the class of functions presented in Proposition 1. This is also required when we estimate models via score minimization, which we will consider in Section 4. Such a dependence of the risk measurement procedure on the choice of scoring functions complicates the fair evaluation of forecasts. Patton (2020) also shows that forecast rankings can be sensitive to the choice of scoring function.
@@
is called an elementary scoring function for VaR. For competing forecasts $\{\widehat{VaR}_{t}^{(\ell)}\}_{t\in\mathbb{T}_{\text{out}}}$, $\ell=1,\dots,L$, of VaRs, a Murphy diagram displays the curves $\eta\mapsto(1/|\mathbb{T}_{\text{out}}|)\sum_{t\in\mathbb{T}_{\text{out}}}\operatorname{S}_{\eta}^{\text{VaR}}\left(\widehat{VaR}_{t}^{(\ell)},s_{t}\right)$, $\ell=1,\dots,L$, against $\eta$. The diagram is negatively oriented in the sense that forecasts with lower curves are evaluated to be more accurate.
The next proposition provides mixture representations of the scoring functions for ESCs presented in Proposition 1.
@@
    The scoring function $\operatorname{S}^{\text{ESC}}$ in Proposition 1 (S2), with $\phi_{j}=:\phi$ for $j=1,\dots,d$, admits the mixture representation:
@@
Remark 2.
Following Section 2.3 of Ehm et al. (2016), the elementary scoring function (6) can be interpreted as a degree of regret for the $j$th branch of a company who has a fixed capital $\eta$ to cover a future loss in distress, whose point forecast is $m_{j}$; see Section S1.3 for details.
Since $\phi_{j}$ and $\phi$ in (M1) and (M2) above are strictly convex, the corresponding measures $H_{j}$ and $H$ assign positive mass to any finite interval on $\mathbb{R}$. Note that $\mathrm{d}H(x)=2\mathrm{d}x$ for the squared loss, and $\operatorname{S}_{j,\eta}^{\text{ESC}}$ arises from $\operatorname{S}_{j}^{\text{ESC}}$ in Proposition 1 (S1) by taking $\phi_{j}(x)=(x-\eta)_{+}$ although this function is not strictly convex.
Based on the mixture representations (5) and (7), a Murphy diagram of ESCs can be drawn analogously to the case of VaR. For $l\in\{1,\dots,L\}$, let $\{\hat{\bm{\varrho}}_{t}^{(\ell)}\}_{t\in\mathbb{T}_{\text{out}}}$, $\hat{\bm{\varrho}}_{t}=(\widehat{ESC}_{1,t}^{(\ell)},\dots,\widehat{ESC}_{d,t}^{(\ell)},\widehat{VaR}_{t}^{(\ell)})$, be a series of predictions in the setting of Section 2.3. The Murphy diagram of the $j$th ESC displays the curves.
@@
The Murphy diagram of the $d$-tuple of ESCs analogously exhibits $\eta\mapsto\bar{\operatorname{S}}_{\eta}^{\text{ESC}}(l):=\sum_{j=1}^{d}\bar{\operatorname{S}}_{j,\eta}^{\text{ESC}}(l)$ for $\ell=1,\dots,L$. To obtain these curves and their differences on the whole real line, it suffices to evaluate them on a finite set because $\bar{\operatorname{S}}_{j,\eta}^{\text{ESC}}(l)$ and $\bar{\operatorname{S}}_{\eta}^{\text{ESC}}(l)$ are piecewise linear functions of $\eta$ with all kinks and jump points contained in $\mathcal{D}_{j}^{(l)}=\{x_{j,t},\widehat{ESC}_{j,t}^{(\ell)},t\in\mathbb{T}_{\text{out}}\}$ and $\mathcal{D}^{(l)}=\bigcup_{j=1}^{d}\mathcal{D}_{j}^{(l)}$, and $\bar{\operatorname{S}}_{j,\eta}^{\text{ESC}}(l)$ and $\bar{\operatorname{S}}_{\eta}^{\text{ESC}}(l)$ vanish outside of $[\min(\mathcal{D}_{j}^{(l)}),\max(\mathcal{D}_{j}^{(l)})]$ and $[\min(\mathcal{D}^{(l)}),\max(\mathcal{D}^{(l)})]$, respectively.
Note that Murphy diagrams of ESCs depend on the corresponding forecasts of total VaR. Consequently, we can conduct a completely fair evaluation among multiple forecasts of ESCs when they share the same forecasts of total VaR. In such a case, more advanced analyses, such as forecast dominance tests proposed by Ziegel et al. (2020), are available.
@@ 4 Proposed models for estimating dynamic ES contributions
In previous sections we have mentioned that clear and rigid comparison among forecasting models of ESCs is feasible among those with common forecasts of total VaRs. In this section we exploit this feature and propose new models based on compositinal regression.
According to the evaluation criterion based on multi-objective elicitability of ESCs, forecast accuracy of total VaR is of utmost importance since estimated ESCs are compared only for models estimating total VaRs with equal accuracy. As considered in Dimitriadis and Hoga (2023), this valuation principle naturally encourages a two-stage approach, where we first model the dynamics of $\{VaR_{t}\}$, possibly induced from the dynamics of $\{S_{t}\}$, and then consider the dynamics of ESCs, which may also be induced from the dynamics of $\{(X_{j},S_{t})\}$ for $j=1,\dots,d$. We summarize benefits and limitations of such a top-down approach in Section S2.1. In this approach the dynamics of $\{ES_{t}\}$ can also be estimated together with $\{VaR_{t}\}$ in the first stage due to the natural order in the risk measurement procedure and, more importantly, joint elicitability of $(\operatorname{VaR}_{\alpha},\operatorname{ES}_{\alpha})$, where ESα can be elicited only in combination with VaRα (Fissler and Ziegel 2016). An alternative approach is to specify the dynamics of $\{\bm{X}_{t}\}$, which we call a bottom-up approach. This approach may be discouraged in terms of the two-stage forecast evaluation of ESCs since total VaR is modeled only indirectly; see Section S2.1 for more details of this approach.
Despite the appeals of the top-down approach in forecasting ESCs, special attention is required in this approach so that the empirical counterpart of FAP $\sum_{j=1}^{d}\widehat{ESC}_{j,t}=\widehat{ES}_{t}$ is satisfied. To handle this constraint, we propose new models for estimating dynamic ESCs. In what follows, we describe our proposed procedure for the case of one-step ahead forecast, where we use the losses $\{\bm{X}_{t}\}_{t\in\mathbb{T}_{\text{in}}}$, $\mathbb{T}_{\text{in}}=\{1,\dots,n\}$, to forecast ESCs at time $n+1$.
The proposed model consists of two stages, where the dynamics of total ES is first estimated in combination with that of total VaR, and then ESCs are estimated from the proportion of the total ES to each component of risk. In the first stage, we estimate the dynamics of $(ES_{t},VaR_{t})$ for $t\in\mathbb{T}_{\text{in}}$ and then forecast $(VaR_{n+1},ES_{n+1})$ based on this estimated dynamics. For this purpose, various models are proposed in the literature; see, for example, Patton et al. (2019)Taylor (2019) and Taylor (2022). Joint elicitability of the pair of risk measures $(\operatorname{VaR}_{\alpha},\operatorname{ES}_{\alpha})$ enables us to estimate the joint dynamics of this pair by score minimization; see, for example, Taylor (2019). In this paper we do not specify any specific model in this stage and instead select the best model based on the currently available information and target models for comparison. Denote by $(\widehat{VaR}_{t},\widehat{ES}_{t})_{t\in\mathbb{T}}$, $\mathbb{T}=\{1,\dots,n+1\}$, the dynamics forecasted in this stage.
In the second stage, we allocate $\widehat{ES}_{t}$ to each component of risk with the weight vector $\bm{w}_{t}=(w_{1,t},\dots,w_{d,t})^{\top}\in\mathbb{R}^{d}$, $\sum_{j=1}^{d}w_{j,t}=1$, such that $ESC_{j,t}=w_{j,t}\times ES_{t}$. We assume that $ESC_{j,t}>0$ for all $j=1,\dots,d$ and $t\in\mathbb{T}$. This assumption implies that $\bm{w}_{t}\in(0,1)^{d}$ by the diversifying property $ESC_{j,t}\leq ES_{t}$ (Kalkbrener 2005). We further discuss this assumption and its relaxation in Remark 3. Consequently, the weight vector $\bm{w}_{t}$ lies in $\mathcal{S}_{d}=\left\{\bm{w}\in(0,1)^{d}:\sum_{j=1}^{d}w_{j}=1\right\}$. Data on $\mathcal{S}_{d}$ are called compositional data, and statistical modeling of such data has been extensively studied in the literature; see Aitchison (1982); Aitchison and J. Egozcue (2005); Pawlowsky-Glahn and Buccianti (2011); and references therein. A common approach for modeling compositional data is first transforming the data to an unconstrained space to eliminate the sum constraint and then constructing a model on the unconstrained space.
One of the most prominent examples of such a transform is the isometric log-ratio transformation (ilr, Egozcue et al. 2003) defined by:
@@
With this choice of $V$, $\operatorname{ilr}_{k}(\bm{w})$, $k\in\{1,\dots,d-1\}$, can be interpreted as a normalized balance between the $k$th weight and the group of the weights over the assets $k+1,\dots,d$. Note that 00 and $1$ cannot be included in the compositional data for the above transformations to be well-defined. This well-known limitation of ilr should not cause a significant problem in our analysis since 00 or $1$ in the allocated weights corresponds to the practically irrelevant case when no capital or full capital is allocated to one asset, respectively.